### v1

# Lin Li1,2∗,  Wei Chen1∗,  Jiahui Li3,  Long Chen1†

Abstract
Recent advances in multi-modal large language models (MLLMs) have significantly improved object-level grounding and region captioning, but remain limited in visual relation understanding (e.g., scene graph generation), particularly in modeling N-ary relationships that identify multiple semantic roles among an action event. Such a lack of semantic dependencies modeling among multi-entities leads to unreliable outputs, intensifying MLLMs’ hallucinations and over-reliance on language priors. To this end, we propose Relation-R1, the first unified relational comprehension framework that explicitly integrates cognitive chain-of-thought (CoT)-guided Supervised Fine-Tuning (SFT) and Group Relative Policy Optimization (GRPO) within a reinforcement learning (RL) paradigm. Specifically, we first establish foundational reasoning capabilities via SFT, enforcing structured outputs with thinking processes. Then, GRPO is utilized to refine these outputs via multi-reward optimization, prioritizing visual-semantic grounding over language-induced biases, thereby improving generalization capability. Extensive experiments on widely-used PSG and SWiG datasets demonstrate that Relation-R1 achieves state-of-the-art performance in both binary and N-ary relation understanding.
11footnotetext: * The first two authors contribute equally to this work.22footnotetext: † Corresponding Author: Long Chen.
1 Introduction
Recent advancements in MLLMs have significantly enhanced holistic image understanding and object-level grounding capabilities, enabling tasks like region captioning and referring question answering. However, current MLLMs exhibit critical limitations in relational scene understanding — a core competency for achieving human-like visual cognition, as illustrated in Figure 1(a).
Figure 1: Illustrations of our Relation-R1 and other MLLMs. (a) Grounded MLLMs: Link objects in text to image regions but fail to capture object relations. (b) Relation Comprehension MLLMs: Model pairwise object relations but lack generalization to N-ary interactions. (c) Unified Relation Comprehension MLLMs: Jointly handle binary and $N$-ary relation detection.
Despite existing studies have made progress in binary relation detection [1, 2], i.e., identifying subject-object interactions (e.g., “child-drinking-glass”), they fail to capture $N$-nary relationship, that identify $N$-th entities engaged in the relation with their roles (what, who, and where concepts). It leads to neglecting the semantic dependencies among multi-entities, such as the glass’s role as the functional container of milk in the drinking action, which may result in unreasonable relation triplets, e.g., “child-drinking-glass”. Furthermore, these unreasonable pair-wise relations may make the MLLMs suffer from both hallucinations and over-relying on language priors rather than visual-semantic evidence. Consequently, there arises an urgent demand for a unified framework capable of addressing these shortcomings. This raises a critical question: How can we devise a unified framework for joint binary and N-ary relational reasoning with grounded cues?
Intuitively, such complicated relational comprehension requires both reasoning and generalization abilities. As for reasoning, the model must perform multi-step and multi-modal reasoning, it includes entity identification, relation recognition, role inference (e.g., “glass” as a container), localization, and contextual integration. In terms of generalization, models trained on limited data struggle to extrapolate to novel entity combinations or ambiguous contexts (e.g., “cup” as a container). In addition, language priors (e.g., common-sense assumptions) often dominate over visual grounding, further limiting their robustness.
Fortunately, DeepSeek-R1-Zero has successfully demonstrated the emergence of reasoning capabilities in LLMs purely through RL. Inspired by this breakthrough, we propose a unified relation comprehension framework Relation-R1, that integrates Supervised Fine-Tuning (SFT) and RL to empower MLLMs with relational reasoning and generalization capability. While SFT alone suffers from poor generalization due to reliance on fixed training patterns, RL struggles to produce structured outputs with consistent formatting [3]. To alleviate these limitations, we design a two-stage pipeline:
Stage 1. SFT: Establishes foundational reasoning through cognitive chain-of-thought guidance. The model generates structured outputs by analyzing visual evidence (e.g., object detection, spatial localization, etc.), ensuring visual grounding over language priors.
Stage 2. RL: Refines outputs via GRPO guided by multi-rewards: (1) format rewards, (2) binary relation rewards, and (3) N-ary relation rewards. This stage employs policy gradient updates to prioritize visual-semantic coherence over language priors, fostering robust relational reasoning and generalization.
By first leveraging SFT to establish reasoning foundations and then applying GRPO to optimize for structured, grounded outputs, Relation-R1 achieves robust relational reasoning and generalization to novel contexts, directly addressing the limitations of prior MLLMs.
We validate the effectiveness of Relation-R1 on both the binary relation detection task (e.g., Scene Graph Generation) and N-ary relation detection task (e.g., Grounded Situation Recognition) under both few-shot and full-data settings. Experiments on PSG [4] dataset and SWiG [5] dataset demonstrate state-of-the-art performance in relational reasoning while maintaining generalization in ambiguous contexts. In summary, our contributions are threefold:
  • We incorporate both binary relation detection and N-ary relation detection in the unified relation comprehension framework.
  • We propose Relation-R1, a unified framework that integrates cognitive CoT guided SFT with RL. To the best of our knowledge, this is the first R1-like work to enable both relational reasoning and robust generalization.
  • Experiments on the prevalent relation understanding datasets under different settings demonstrate the effectiveness of Relation-R1.
2 Related Work
Reasoning in Multimodal Large Language Models. Recent Large Language Models (LLMs) have demonstrated advanced reasoning capabilities by emulating human-like stepwise thinking, significantly enhancing performance on complex tasks [6, 7]. A pivotal breakthrough, DeepSeek-R1 [7], leverages large-scale reinforcement learning (RL) with formatting and result-oriented rewards to enable LLMs to autonomously generate human-like complex Chain-of-Thought (CoT) reasoning, achieving state-of-the-art results across diverse domains. Multimodal Large Language Models (MLLMs) extend these capabilities through systematic integration of cross-modal knowledge representation and task-specific rewards [8, 9, 10] For multimodal mathematical reasoning, frameworks such as R1-Onevision [8], Virgo [9], and MM-Eureka [11] jointly reason over visual and textual inputs to solve quantitative tasks, while Vision-R1 [10] combines cold-start initialization with GRPO to refine complex CoT reasoning. For fine-grained classification and grounding tasks, Visual-RFT [12] and R1-Omni [13] utilize GRPO-based reinforcement algorithms and verifiable rewards to enhance contextual reasoning. As for the pixel-level understanding task, SegZero [14] employs reinforcement learning to achieve high-resolution semantic segmentation. Despite these advances, existing MLLMs predominantly focus on object-centric recognition rather than unified visual relation understanding, which requires compositional reasoning about geometric, semantic, and functional interactions between objects and contexts. To bridge this gap, our Relation-R1 explicitly integrates relational reasoning into MLLMs via reinforcement learning, enabling systematic inference for both scene interpretation and event prediction.
Scene Graph Generation (SGG). SGG is a critical task in scene understanding, with prior approaches falling into two categories: 1) Two-stage SGG, which sequentially detects objects and infers pairwise relations but suffers from error propagation [15, 16, 17, 18]; 2) One-stage SGG, which unifies detection and relation prediction in end-to-end frameworks (e.g., DETR-based methods [19, 20, 21]) but often lacks multi-role involved relation modeling. Recent efforts in open-vocabulary SGG (OVSGG) leverage Vision-Language Models (VLMs) [22, 23, 24] or MLLMs [1, 2, 25] to handle novel entities/relations, yet limited modeling of complex multi-entity interactions, and suffer from overfitting in base categories during SFT. Our Relation-R1 addresses OVSGG challenges by integrating a cognitive CoT-guided RL framework to model complex multi-entity interactions and mitigate SFT overfitting, enabling robust zero-shot reasoning without predefined category constraints.
Grounded Situation Recognition (GSR). GSR extends scene understanding by jointly modeling actions (verbs) and their associated semantic roles. Pratt et al. [5] pioneered this field with a two-stage RNN-based framework: verb detection followed by noun localization. Subsequent works [26, 27, 28, 29] improved this pipeline by integrating transformers and semantic relation modeling, enabling more coherent scene interpretations. Despite these efforts, previous efforts that depend on annotated training samples may face challenges like long-tailed distribution [15, 16, 17], potentially limiting real-world applicability. While LEX [30] realizes training-free zero-shot GSR, it fails to provide an end-to-end framework and exhibits suboptimal performance and computational efficiency due to its reliance on LLMs’ generated descriptions. This paper proposes Relation-R, a unified framework that bridges the gap between open-ended capabilities and end-to-end training.
3 Methodology
Figure 2: Overview of the proposed Relation-R1 framework. (a) SFT with cognitive CoT, which enforces format alignment through task-specific reasoning processes. (b) RL with GRPO, employs rule-based rewards (e.g., format rewards, binary relation rewards, and N-nary relation rewards) to cultivate robust relation reasoning and generalization ability through policy gradient updates.
3.1 Preliminaries
Binary Relation Detection. This task requires the model to generate an open-ended scene graph given an image, explicitly marking semantic relationships between pairwise objects [1]. The output is structured as a sequence of relation triplets annotated with formal tags: objects are denoted by <ref> tags followed by their bounding box coordinates < box> to localize the entity in the image, e.g., <ref>person</red></box>[[$x_{1}$, $y_{1}$, $y_{x}$, $y_{2}$]]</box>. Predicates are enclosed in <pred> tags, followed by references to their subject and object entities via their corresponding bounding boxes.
N-ary Relation Detection. Similarly, this task aims to output a open-ended grounded situation frame [5] structured as a primary activity (relation), followed by a sequence of <role> tags along with corresponding bounding box coordinates, each specifying an entity’s role, e.g., <agent>person</agent></box>[$x_{1}$, $y_{1}$, $y_{x}$, $y_{2}$]</box>. Each role explicitly defines the entity’s participation in the activity.
3.2 Supervised Fine-Tuning
As mentioned above, directly adopting GRPO faces challenges in generating format-inconsistent scene graph captions and grounded situation frame outputs. To this end, we adopt supervised fine-tuning that serves as a foundational stage for task-specific format alignment (c.f., Figure 2(a)).
However, utilizing merely the answer format as the SFT target risks compromising the model’s reasoning capability by encouraging rigid adherence to syntactic structures [3]. To mitigate this, we devise the cognitive Chain-of-Thoughts (CoT), which incorporates stepwise cognitive processes such as object classification, object grounding, visual relationship inference, etc. This ensures the model retains its ability to perform multi-step reasoning while adhering to output formats.
To avoid the MLLM overfitting to specific CoT patterns during SFT, we devise a cognitive CoT generation prompt that explicitly guides an extra strong MLLM (e.g., Qwen 2.5-VL [31]) to generate step-by-step reasoning processes. To be specific, this cognitive CoT generation prompt consists of task definition, scene graph caption/grounded situation frame examples, and GT scene graph caption/grounded situation frame of the current image. The former two make the MLLM generate the correct task format. Meanwhile, the latter one anchors reasoning content to visual evidence, ensuring that the MLLM’s cognitive CoT aligns with both semantic validity and visual grounding rather than relying on category priors. The detailed Cognitive CoT generation prompt is left in the Appendix.
These cognitive CoT are explicitly encapsulated within <think> tags, during SFT. By leveraging this structured training signal, the MLLM learns to generate format-compliant outputs and grounded reasoning processes as an initial capability, which is critical for subsequent CoT-guided generalization under the GRPO.
3.3 Group Relative Policy Optimization
Following the Deepseek-R1 [7], the Group Relative Policy Optimization (GRPO) is adopted as the reinforcement learning (RL) algorithm for optimizing our Relation-R1 framework, as illustrated in Figure 2(b). Distinct from critic-dependent methods [32], GRPO directly calculates advantage estimates through response group comparisons sampled from the policy model, inherently eliminating the need for an extra critic network and reducing computational complexity [7]. Key components include the a question $q$, policy model $\pi_{\theta}$, response group $\{o_{i}\}$, and a frozen reference model $\pi_{\theta_{\text{ref}}}$, which acts as a regularization baseline to preserve prior knowledge. The GRPO objective function $J_{\text{GRPO}}(\theta)$ balances reward maximization and policy stability by leveraging following objectives:
$J_{\text{GRPO}}(\theta)$ $=\mathbb{E}_{q\sim\mathcal{Q},\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}}$ (1)
$\left[\frac{1}{G}\sum_{i=1}^{G}\min\left(\rho_{i}A_{i},\text{clip}(\rho_{i},1-\epsilon,1+\epsilon)A_{i}\right)-\beta D_{\text{KL}}(\pi_{\theta}\|\pi_{\text{ref}})\right],$
where $\rho_{i}$ = $\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{\text{old}}}(o_{i}|q)}$ quantifies policy change, $\epsilon$ controls clipping thresholds, and $\beta$ penalizes deviations via the KL divergence term. The advantage score $A_{i}$ is standardized as:
$A_{i}=\frac{r_{i}-\text{mean}\left(\{r_{1},\ldots,r_{G}\}\right)}{\text{std}\left(\{r_{1},\ldots,r_{G}\}\right)},$ (2)
here, $r_{i}$ denotes the reward for response $o_{i}$. The KL divergence term enforces proximity to the reference policy is formulated as:
$D_{KL}(\pi_{\theta}\|\pi_{\text{ref}})=\frac{\pi_{\text{ref}}(o_{i}|q)}{\pi_{\theta}(o_{i}|q)}-\log\left(\frac{\pi_{\text{ref}}(o_{i}|q)}{\pi_{\theta}(o_{i}|q)}\right)-1,$ (3)
which ensures controlled exploration without excessive divergence from $\pi_{\theta_{\text{ref}}}$.
The reward function $r_{i}$ is formulated to align with the model’s learning objective for the unified relation understanding. It incorporates rule-based rewards, which consist of three crucial components: Format Compliance Rewards, Binary Relation Rewards, and $N$-nary Relation Rewards. The former enforces syntactic validity by penalizing outputs that deviate from predefined structural constraints. The latter two components focus on task-specific performance.
Format Rewards. This reward enforces strict adherence to the reasoning template format: <think> thinking process </think><answer> answer </answer>. The reward function is defined as:
$r_{\text{form}}(o_{i})=\begin{cases}1&\text{if }o_{i}\text{ adheres to the format},\\ 0&\text{otherwise}.\end{cases}$ (4)
This binary mechanism ensures syntactic compliance while allowing content flexibility.
Binary Relation Rewards. Inspired by established evaluation practices [4, 1], we devise the sample-level triplet Recall and sample-level mean Recall (mRecall) for each predicate category in the response of each sample. Specifically, a triplet is deemed correct if:
  • Triplet Accuracy: All entity categories (subject, predicate, object) in this sample are correctly identified.
  • Spatial Localization: The predicted bounding boxes for the subject and object achieve an Intersection over Union (IoU) >= 0.5 with their respective ground-truth boxes.
The final binary relation reward is computed as a weighted combination of these two metrics:
$r_{\text{binary}}(o)=\alpha\cdot Recall+(1-\alpha)\cdot mRecall,$ (5)
where $\alpha$ balances the trade-off between two metrics.
N-ary Relation Rewards. Similarly, to evaluate higher-order relational structures in grounded situation frames, we introduce N-ary Relation Rewards, which assess the model’s ability to detect primary activity and multi-entity interactions. Specifically, we calculate the verb accuracy $Acc_{v}$ and noun accuracy $Acc_{n}$:
  • Verb Accuracy: The predicted primary activity (verb) in the grounded situation frame is the same as the ground-truth activity.
  • Noun Accuracy: All entities involved in the event must satisfy IoU >=0.5 with their ground-truth bounding boxes, and their roles must align with the event’s semantic structure.
The eventual N-ary relation reward is computed as a weighted combination of these two accuracies:
$r_{\text{n-ary}}(o)=\beta\cdot Acc_{v}+(1-\beta)\cdot Acc_{n},$ (6)
where hyperparameter $\beta$ balances the trade-off between event type precision and multi-entity structural accuracy. During training, we design a multi-task gate that dynamically selects task-specific rewards based on the presence of <ref> tags in the solutions.
4 Experiment
4.1 Experimental Setup
4.1.1 Datasets
PSG. The Panoptic Scene Graph (PSG) dataset [4] is built upon the COCO dataset [33], comprising 48,749 annotated images of which 46,563 are for training and 2,186 are for testing. It integrates 80 “thing” object categories and 53 “stuff” object categories, alongside 56 relation categories. To enable open-ended scene graph generation, we followed ASMv2 [1], using scene graph caption where objects are tagged with <ref> and paired with bounding box coordinates specifying their locations, while predicates are marked with <pred> and linked to two bounding boxes indicating their subject and object entities for training.
SWiG. The SWiG dataset [5] extends the imSitu dataset [34] by adding extra bounding box annotations for visible semantic roles, with 69.3% of semantic roles localized. Each image is annotated with a verb and a variable number of semantic roles (1$\sim$6 roles per image), where each verb is associated with three verb frames annotated by independent human annotators. The dataset comprises 25,200 testing images spanning 504 verb categories, 190 semantic role categories, and 9,929 noun entity categories. Unlike traditional output structures that merely predict categories and bounding boxes, we constructed grounded situation frames by filling verb-centric templates (e.g., “The AGENT drinks a LIQUID from a CONTAINER”) with three critical components: role tags (e.g., <agent>, entity categories (e.g., child), and coordinates of corresponding bounding boxes in <box> tags.
4.1.2 Evaluation Metrics
Binary Relation Detection. Following the standard SGG evaluation protocol [15, 1], we reported 1) Recall: The ratio of correctly predicted triplets to the total number of ground-truth triplets, requiring both label accuracy and bounding box IoU >= 0.5 for subject and object. 2) mean Recall (mRecall): The average recall across all predicate categories.
N-ary Relation Detection. We adopted the same five evaluation metrics as [5]: 1) verb: the accuracy of verb prediction. 2) value: the noun accuracy for individual semantic role. 3) value-all (val-all): Overall noun prediction accuracy across all semantic roles in an event. 4) grounded-value (grnd): Bounding box accuracy for each semantic role, requiring an IoU >= 0.5 between predicted and ground-truth boxes. 5) grounded-value-all (grnd-all): entire bounding box accuracy across all semantic roles in an event.
4.1.3 Implementation Details
We utilized Qwen2.5-VL-3B [31] as our base model. Relation-R1 was trained on an 8xA100 GPU server. During the SFT phase, we employed a total batch size of 64, with 8 samples processed per training step. For the GRPO component, we used a total batch size of 32. The learning rates were set to 1e-5 for SFT and 1e-7 for GRPO, respectively.
Table 1: Performance(%) comparison with various SGG models on the PSG dataset [4].
            Method Model Size Answer Recall mRecall Mean
IMP [35] ${}_{\text{CVPR'17}}$ Close-ended 16.50 6.50 11.50
MOTIFS [16] ${}_{\text{CVPR'18}}$ 20.00 9.10 14.55
VCTree [36] ${}_{\text{CVPR'19}}$ 20.60 9.70 15.15
GPSNet [37] ${}_{\text{CVPR'20}}$ 17.80 7.00 12.4
PSGFormer [4] ${}_{\text{ECCV'22}}$ 18.60 16.70 17.65
TextPSG [38] ${}_{\text{ICCV'23}}$ Open-ended 4.80 - -
ASMv2 [1] ${}_{\text{ECCV'24}}$ 13B 14.20 10.30 12.23
LLaVA-SpaceSGG [2] ${}_{\text{WACV'25}}$ 13B 15.43 13.23 14.33
Relation-R1 (Ours) 3B 19.59 18.30 18.95
Table 2: Performance(%) comparison with various GSR models on the SWiG [5] dataset.
         Method Answer Verb Value Value-all Gnd Value Gnd Value-all
ISL [5] ${}_{\text{ECCV'20}}$ Close-ended 39.36 30.09 18.62 22.73 7.72
JSL [5] ${}_{\text{ECCV'20}}$ 39.94 31.44 18.87 24.86 9.66
GSRTR [26] ${}_{\text{BMVC'21}}$ 40.63 32.15 19.28 25.49 10.10
CoFormer [28] ${}_{\text{CVPR'22}}$ 44.66 35.98 22.22 29.05 12.21
SituFormer [27] ${}_{\text{AAAI'22}}$ 44.20 35.24 21.86 29.22 13.41
GSRFormer [29] ACM ${}_{\text{ACM MM'22}}$ 46.53 37 .48 23.32 31.53 14.23
OpenSU [39] ${}_{\text{ICCVW'23}}$ Open-ended 50.10 41.20 26.56 34.27 15.70
Relation-R1 (Ours) 53.76 43.55 28.23 36.79 26.84
Table 3: Few-Shot GSR performance comparison. $\dagger$ denotes employing the SFT phrase without CoT.
         Method Setting Verb Value Value-all Gnd Value Gnd Value-all
GSRTR [26] ${}_{\text{BMVC'21}}$ One-Shot 0.35 0.10 0.01 0.08 0.00
CoFormer [28] ${}_{\text{CVPR'22}}$ 0.63 0.23 0.02 0.14 0.01
Relation-R1 (Ours) 27.31 10.21 1.43 9.35 1.22
Relation-R1 (Ours) 31.67 15.37 4.15 14.01 3.41
GSRTR [26] ${}_{\text{BMVC'21}}$ 4-Shot 6.95 4. 32 1.83 2.79 0.65
CoFormer [28] ${}_{\text{CVPR'22}}$ 12.39 8.27 3.93 5.28 1.25
Relation-R1 (Ours) 34.88 17.14 3.94 15.63 3.38
Relation-R1 (Ours) 38.02 21.87 7.39 19.64 5.92
(a) SGG.
(b) GSR.
Figure 3: The statistics of rewards during training.
Figure 4: Performance(%) comparison between ASMv2 and Relation-R1 over all predicates on test set of PSG. The orange area denotes the predicate distribution of the training set.
Figure 5: Qualitative results of binary relation detection (scene graph generation).
Figure 6: Qualitative results of N-ary relation detection (grounded situation recognition).
4.2 Quantitative Comparison Result
4.2.1 Binary Relation Detection
Open-ended SGG. We evaluated Relation-R1 on PSG, comparing it against various established and recent SGG models, as detailed in Table 1. Positioned within the open-ended generation category, Relation-R1 achieved SOTA performance with the highest mRecall of 18.30% and overall Mean of 18.95%, substantially surpassing prior open-ended methods like LLaVA-SpaceSGG (13.23% mRecall, 14.33% Mean). It also achieved a competitive Recall of 19.59%. Notably, Relation-R1 attained these leading results using a significantly smaller 3B parameter model compared to the 13B models used by recent competitors like ASMv2 and LLaVA-SpaceSGG. In addition, we visualized predicate-level comparison against ASMv2 and the predicate distribution in Figure 4. As seen, Relation-R1 outperforms ASMv2 by a clear margin on both prevalent and rare predicates. These results clearly validated the effectiveness and efficiency of our approach for the SGG task in the open-ended setting.
4.2.2 N-ary Relation Detection
Open-ended GSR. We evaluated Relation-R1 for N-ary relation detection on SWiG, comparing it against prior work in Table 2. Within the open-ended setting, Relation-R1 established a new SOTA, achieving top performance across all reported metrics: Verb (53.76%), Value (43.55%), Value-all (28.23%), Gnd Value (36.79%), and Gnd Value-all (26.84%). Notably, Relation-R1 significantly surpassed the previous best open-ended model, OpenSU [39], exhibiting substantial gains, particularly with a remarkable +11.14% absolute improvement on the challenging Gnd Value-all metric. This clearly demonstrated Relation-R1’s superior capacity for comprehending and grounding complex N-ary relationships in visual scenes.
Few-Shot GSR. We also evaluated Relation-R1’s few-shot generalization capability (Table 3). In the one-shot setting, Relation-R1 achieved dramatic improvements over prior work, reaching 31.67% (Verb) and 3.41% (Gnd Value-all) compared to CoFormer’s [28] 0.63% and 0.01%, respectively. This advantage is maintained in the 4-shot setting, where Relation-R1 reached 38.02% (Verb) and 5.92% (Gnd Value-all), significantly surpassing previous methods. These results clearly demonstrate Relation-R1’s superior ability to generalize and perform robust N-ary relation detection, setting a new SOTA for few-shot GSR.
4.3 Qualitative Result
Figure 5 and Figure 6 provided qualitative illustrations of Relation-R1. For Binary Relation Detection (Figure 5), the visualized thinking process reveals how Relation-R1 effectively decomposes the complex task into constituent cognitive steps: object classification, localization, and relation inference. For instance, after identifying the central bed, Relation-R1 pinpoints related objects like the rug and chair and explicitly grounds the inferred spatial relationships (on, beside) between them. This systematic approach demonstrates the successful decomposition of the scene into accurate, grounded object-relation pairs. Similarly, for N-ary Relation Detection (Figure 6), Relation-R1 showcases its understanding by first identifying the core action (crowning) and then methodically grounding each associated semantic role – Agent (woman), Object (tiara), and Recipient (woman) – to its corresponding entity and bounding box. This structured analysis underscores Relation-R1’s ability to parse complex events involving multiple participants and their distinct roles within the scene.
4.4 Diagnostic Experiment
To gain more insights into Relation-R1, we conducted a set of comprehensive studies on SGG and GSR.
Reward Statistics. We visualized the variation curves of rewards in the SGG task and the GSR task, respectively, as displayed in Figure 3(a) and Figure 3(b). It can be seen that at the beginning stage, both tasks can obtain decent rewards due to the foundation laid by SFT. With the gradual progress of GRPO, the rewards of both tasks show a gradually increasing trend.
5 Conclusion
This paper reveals critical limitations of existing MLLMs in visual relation understanding, particularly concerning N-ary relationships and the detrimental effects of language-induced bias. We propose Relation-R1, an advanced, unified framework designed to jointly tackle both binary and N-ary relational comprehension. Our core innovation lies in a two-stage approach: first establishing foundational reasoning and structured outputs via cognitive CoT-guided SFT, and subsequently refining the model’s outputs using RL with GRPO. This RL stage, guided by carefully designed multi-rewards, explicitly promotes visual-semantic grounding and robust generalization. Extensive experiments conducted on the challenging PSG and SWiG datasets validate the effectiveness of Relation-R1, demonstrating SOTA performance across both binary and N-ary relation detection tasks. By successfully integrating CoT-guided SFT with R1-inspired RL refinement, Relation-R1 represents a solid step towards equipping MLLMs with more human-like relational reasoning capabilities.